OpenStudy (anonymous):

Is 6/-4 the slope for the points 1,-4 and -3,2

OpenStudy (anonymous):

yup

OpenStudy (whpalmer4):

slope is difference in the y values divided by the difference (in the same order!) of the x values if we have \((1,-4)\) as the first point and \((-3,2)\) as the second point, \[m = \frac{-4-2}{1-(-3)} = \frac{-6}{4} = \frac{6}{-4} = -\frac{6}{4}\]

OpenStudy (anonymous):

Reduce to - 3/2